# Accounting for multiplicity in machine learning benchmark performance

### v5

@@
Accounting for multiplicity in machine learning benchmark performance
Kajsa Møllersen1 and Einar Holsbø2
(1Department of Community Medicine
2Department of Computer Science
UiT - The Arctic University of Norway
@@ 1 Estimating state-of-the-art performance
In this article, we direct attention to the reporting of state-of-the-art classifier performance by presenting the sample maximum estimator distribution function and analysing its bias. This analysis is done using the framework of multiplicity.
Within the fields of research that develop, refine, test and use classifiers, the term state-of-the-art (SOTA) commonly refers to the highest reported performance to date, and is used as a reference point when presenting and comparing models. The reported performance of a classifier, originating from applying that classifier to a specific data set, a sample, is in effect an estimate of that classifier’s underlying population performance.
The highest reported performance among a number of classifiers serves as an estimate of SOTA, although the term ’estimate’ is seldom used explicitly in the SOTA context. The highest reported performance corresponds to the sample maximum of population performances, and this estimator is biased when used to estimate SOTA. The sample maximum gives in overoptimistic estimates of the maximum population performance, and is problematic as a reference point in research. The mechanisms creating this bias comes from multiplicity, and the fact that the we pick the highest sample performance among multiple classifier results.
In its simplest form, with independent binary classifiers with identical probability of success, the distribution function of the sample maximum estimator exists in closed form with only a few parameters, and can be directly applied to calculate and analyse the bias in a certain situation. With growing complexity, such as mutual dependencies and more elaborate performance measures, the distribution function of the estimator is no longer known, but the mechanisms at play and the pitfalls to be aware of are much the same, and limited studies can be done by straightforward simulations.
@@
Machine learning models are commonly evaluated and compared by their performance on data sets from public repositories, where models can be evaluated under identical conditions and across time. Theoretical results play an important role in machine learning research, but for the last couple of decades the success of increasingly complex models has pushed the proof towards the pudding: performance on benchmark data sets.
The numbers don’t lie: if model A performs better than model B on a specific data set, that is what it did. But the truth that these numbers tell is limited. We are interested in the population performance, and the sample performance is only an estimate. A good estimating procedure gives a reliable result, and there are plenty of pitfalls to be aware of: Overfitting to the data set, originating from wrong use of cross-validation, data leakage and data reuse, can give overly optimistic estimates. An independent test set, often referred to as a hold-out set, is a sample that is not used for development of the model, and prevents overfitting. Competitions and challenges with prize money typically withhold the test set so that it becomes impossible for the participants to use it for training.
The highest ranked performance on a problem is referred to as SOTA performance, and is used, among other things, as a reference point for publication of new models. As a specific model’s performance on a test set is just an estimate of its true underlying performance, the reported SOTA performances are estimates, and hence come with bias and variance. The practice of hold-out test sets in competitions safe-guards against overoptimistic estimates for the individual classifiers. Data is often released after the competition has ended, and new methods can be compared to the SOTA performance.
As we will show in the subsequent sections, using the highest-ranked performance as an estimate for SOTA is a biased estimator, giving overoptimistic results. The mechanisms at play are those of multiplicity, a topic that is well-studied in the context of multiple comparisons and multiple testing, but has, as far as the authors are aware of, been nearly absent from the discussion regarding SOTA estimates.
Biased SOTA estimation has consequences both for practical implementations and for research itself. A health care provider might want to use some machine learning method for tumour detection, but only if the method correctly classifies at least 80%. They follow updates on SOTA performance, and are ready to invest time and money in a pilot project once the lower bound surpasses 80%. If the SOTA performance is an optimistic estimate, the health care provider has now lost time and money, and perhaps also some faith in research results.
For the research community, the situation is perhaps even more dramatic. The optimistic SOTA estimate is used as a standard for evaluating new methods. Not to say that a new method necessarily needs to beat the current SOTA to catch interest from reviewers and editors, but it certainly helps if its performance is comparable. It has been pointed out (see Section 2) that the incremental increase in performance can obstruct the spread of new ideas that have substantial potential, but just haven’t reached their peak yet.
In this article, we provide a probability distribution for the case of multiple classifiers. We demonstrate the impact of multiplicity through simulated examples. We show how classifier dependency impacts the variance, but also that the impact is limited when the probability of success is high. Finally, we discuss three real-world examples that each highlights different aspects multiplicity and SOTA estimation. All code is available at https://github.com/3inar/ninety-nine.
@@ 2 Related work
Multiple comparisons is a well-studied subject within statistics, see, e.g., Tukey (1991) for an interesting discussion on some topics of multiple comparisons, among them the confidence intervals (CIs). For a direct demonstration of how multiple testing can lead to false discoveries, Bennett et al. (2009) give an example where brain activity is detected in (dead) salmon “looking” at photos of people. In the machine learning context, with multiple data sets and classifiers, Demšar (2006) suggests adjusting for multiple testing, according to well-known statistical theory. Our article focuses on the estimate itself and its uncertainty, and not comparison of methods.
@@
The re-use of benchmark datasets has been addressed in several publications. Salzberg (1997) point to the problem that arises because these datasets are static, and therefore run the risk of being exhausted by overfitting. Exhausted public data sets still serve important functions, e.g., a new idea can show to perform satisfactory, even when it is not superior. Thompson et al. (2020) distinguish between simultaneous and sequential multiple correction procedures. In sequential testing, which is the reality of public datasets, the number of tests is constantly updated. An interesting topic they bring up is the conflict between sharing incentive, open access and stable false positive rate. There are several demonstrations of the vulnerability of re-use of benchmark datasets. Teney et al. (2020) performed a successful attack on the VQA-CP benchmark dataset, that is, they are able to attain high performance with a non-generalisable classifier. They point towards the risk of ML methods capturing idiosyncrasies of a dataset. Torralba and Efros (2011) introduced the game Name That Dataset! motivating their concern of methods being tailored to a dataset rather than a problem. The lack of statistically significant differences among top performing methods is demonstrated by, e.g., Everingham et al. (2015) and Fernández-Delgado et al. (2014).
@@
The same mechanisms of overfitting are present in challenges through multiple submissions to the validation set. Blum and Hardt (2015) introduced an algorithm that releases the public leaderboard score only if there has been significant improvement compared to the previous submission. This will prevent overfitting by hindering adjustments to random fluctuations in the score and not true improvements of the method. Dwork et al. (2017) used the principles of differential privacy to prevent overfitting to the validation set.
@@
Several publications show that the overfitting is not as bad as expected, and provide possible explanations. Recht et al. (2019) constructed new, independent test sets of CIFAR-10 and ImageNet. They found that the accuracy drop is dramatic, around 10%. However, the order of the best performing methods is preserved, which can be explained by the fact that multiple classes (Feldman et al. (2019)) and model similarity (Mania et al. (2019)) slow down overfitting. Roelofs et al. (2019) compared the results from the private leaderboard (one submission per team) and the public leaderboard (several submissions per team) in Kaggle challenges. Several of their analyses indicate overfitting even though their main conclusion is that there is little evidence of overfitting.
@@
Overfitting and its consequences are avoided in most challenges by withholding the test set so that a method can only be evaluated once and not adapted to the test set performance. Hold-out datasets also prevents wrong use of cross-validation, see, e.g., Fernández-Delgado et al. (2014) and the seminal paper of Dietterich (1998). Another contribution from public challenges is the requirement of reproducibility to collect the prize money, which is a well-known concern in science, see, e.g., Gundersen and Kjensmo (2018). Public challenges typically rank the participants by a (single) performance measure, and can fall victims of Goodhart’s law, paraphrased as “When a measure becomes a target, it ceases to be a good measure.” by Teney et al. (2020). Ma et al. (2021) offered a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
@@
The robustness/stability of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges does not offer an immediate solution to. Bouthillier et al. (2021) showed empirically that hyperparameter choice and the random nature of the data are two large contributors to variance, whereas data augmentation, dropouts, and weights initialisations are minor contributors. Choice of hyperparameters are often adjusted to validation set results, and the strategies of Blum and Hardt (2015) and Dwork et al. (2017) can hinder that, if implemented by the challenge hosts. Our contribution is a focus on better estimate for SOTA performance under the current conditions, where the approach of Bouthillier et al. (2021) does not apply. Bousquet and Elisseeff (2002) investigated how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
@@
There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, e.g., Varoquaux and Cheplygina (2022) for a much broader overview than the one given here. Concerns and solutions regarding multiplicity and better performance estimates are active research fields. We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of SOTA performance.
@@ 3 Multiple classifiers and biased state-of-the-art estimation
3.1 Notation
@@
$\,\,\,\,X\sim\mathcal{B}(n,1\text{-}\theta)$
$\hat{\theta}(X)=\frac{n-X}{n}$ $:\text{estimator for }\theta$
$\hat{\theta}=\frac{n-x}{n}$ $:\text{estimate of }\theta$
$P(X=x)$ $:\text{probability of $x$ failures in one experiment}$
$P_{x}=P(X\leq x)$ $:\text{probability of at most $x$ failures in one experiment}$
$\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ $:\text{number of failures for classifiers }j=1,\ldots,m$
$C_{x}$ $:\text{number of classifiers with at most $x$ failures}$
@@
[…]
$\,\,\,\,z=0,1,2,\ldots,n$

### v6

@@
Abstract
State-of-the-art (SOTA) performance refers to the highest performance achieved by some model on a test sample, preferably under controlled conditions such as public data (reproducibility) or public challenges (independent sample). Thousands of classifiers are applied, and the highest performance becomes the new reference point for a particular problem. In effect, this set-up is an estimate of the expected best performance among all classifiers applied to a random sample; a sample maximum estimate. In this paper, we argue that SOTA should instead be estimated by the expected performance of the best classifier, which can be done without knowing which classifier it is.
Our contribution is the formal distinction between the two, and an investigation into the practical consequences of using the former to estimate the latter. This is done by presenting sample maximum estimator distributions for non-identical and dependent classifiers. We illustrate the impact on real world examples from public challenges.
@@ 1 Estimating state-of-the-art performance
Machine learning (ML) models are commonly evaluated and compared by their performance on data sets from public repositories, where models can be evaluated under identical conditions and across time. Theoretical results play an important role in ML research, but for the last couple of decades the success of increasingly complex models has pushed the evidence towards performance on benchmark data sets. In this paper, we direct attention to the reporting of state-of-the-art (SOTA) performance. We restrict our attention to classifier performance, but the overall mechanisms apply to other models. The term performance will in general refer to quantification of the classifier’s ability to correctly classify the instances in a sample, and we will specify when necessary. Other performance measures, such as computation time, will not be addressed here.
Because of the complexity of both models and their training and test samples, the performances are estimated by applying the models to data samples. Within the fields of research that develop, refine, test and use classifiers, the term SOTA commonly refers to the highest reported performance to date, and is used as a reference point when presenting and comparing models. The highest performance serves as an estimate of SOTA, although the term “estimate” is seldom used explicitly. There is an important distinction between best performance among a number of classifiers, and performance of best classifier, and this paper makes an effort to bring attention to this distinction.
1.1 SOTA performance
SOTA refers to the best model, and hence SOTA performance is the performance of that model. Let $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$ be the probability of success for $m$ classifiers, and let $\hat{\theta}_{j}$, $j=1,\ldots,m$, be their corresponding unbiased estimates. In a public challenge, $\Theta$ would be the classifier performances of the $m$ participants. Taking into account that the participants of a challenge is somewhat arbitrary, and that their classifiers are adjusted to a training sample, we model the classifier ensemble as random sample where $\theta_{j}\sim P(\theta)$, and define SOTA performance as follows
Definition 1 (SOTA classifier performance) Let $\theta_{j}$ be the random variable that denotes the performance of classifier $j$. In an ensemble of $m$ classifiers with performances $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$, the SOTA classifier performance is the expected best performance
$\theta_{SOTA}=\text{E}\max\Theta.$ (1)
For non-random classifier ensembles, the definition holds since the expectation of a constant is just the constant itself, $\theta_{SOTA}=\max\Theta$. This definition answers the question what can we expect the best classifier’s performance to be?, when applied to a new data sample. Note that picking the highest estimated performance derived from an ensemble of classifiers applied to a test sample does not answer this question.
1.2 Sample maximum performance
ML models are commonly ranked by their performance on benchmark datasets, published as leaderboards, often in public challenges where the test sample is withheld to avoid overfitting. The best performance to date, or among the participants in a challenge, is referred to as SOTA, see, for example paperswithcode.com. The published performances are estimates, $\hat{\theta}_{j}$’s, based on the application of models on a specific benchmark or test sample. Although seldom formalised, these are often unbiased estimates, and we will treat them as such. The test samples act as random samples, all the while the point of reporting these performances is the assumption that they can be generalised to similar data sets. Hence, the $\hat{\theta}_{j}$’s are a random sample, and the best performance is the sample maximum estimate, which we define for clarity.
Definition 2 (Sample maximum estimator) Let $\theta_{j}$ be the random variable that denotes the performance of classifier $j$ and let $\hat{\theta}_{j}(X)$ be the corresponding unbiased estimator. For an ensemble of $m$ classifiers with performances $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$, the sample maximum performance estimator is
$\operatorname{\hat{\theta}_{\max}(\mathbf{X})}=\text{E}\max(\hat{\Theta}(\mathbf{X})),$ (2)
where $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ quantifies successes for classifiers $j=1,\ldots,m$.
The sample maximum answers the question What is the best expected performance of an ensemble of classifiers? When $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$ is used as an estimator for $\theta_{SOTA}$, it is biased. The two terms look similar at a glance, the only thing distinguishing them are $\hat{\Theta}(\mathbf{X})$ vs $\Theta$. Taking into account that $\hat{\theta}(X)$ is an unbiased estimator for $\theta$, the distinction can easily be overlooked. The difference reflects a multiplicity effect.
1.3 Multiplicity effect
Multiplicity is, simply put, how the probability of an outcome changes when an experiment is performed multiple times. As we will show in the subsequent sections, using the maximum performance of an ensemble as an estimate for SOTA is a biased estimator giving overoptimistic results. The bias is a multiplicity effect; multiple models being applied to the same sample.
The optimistic estimate is often used as a reference point for evaluating new models. Not to say that a new model necessarily needs to beat the current SOTA to catch interest from reviewers and editors, but it certainly helps if its performance is comparable. Many aspects of public benchmark datasets and public challenges have been thoroughly discussed, and we present some of the topics connected to multiplicity in Section 2. As far as the authors are aware of, the multiplicity effect has been nearly absent from the discussion regarding SOTA estimates.
We present the sample maximum estimator distribution function and its bias for some classifier scenarios. In its simplest form, independent classifiers with identical probability of success, the distribution function of the sample maximum estimator exists in closed form. With growing complexity, such as mutual dependencies and more elaborate performance measures, the distribution function of the estimator is no longer known, but the mechanisms at play and the pitfalls to be aware of are much the same, and limited studies can be done by straightforward simulations. We demonstrate the impact of multiplicity through simulated examples, and show how classifier dependency impacts bias and variance.
Finally, we discuss three real-world examples that each highlights different aspects of multiplicity and SOTA estimation. From these real-world examples, we suggest a SOTA estimate, in line with Definition 1. This is a new perspective for multiplicity effects in the ML context, a step aside from overfitting and re-use of test sets.
All code is available at https://github.com/3inar/ninety-nine.
@@ 2 Related work
Multiple comparisons is a well-studied subject within statistics, see, for example, Tukey (1991) for an interesting discussion on some topics, among them confidence intervals (CIs). For a direct demonstration of how multiple testing can lead to false discoveries, Bennett et al. (2009) gave an example where brain activity is detected in (dead) salmon “looking” at photos of people. In the ML context, with multiple data sets and classifiers, Demšar (2006) suggested adjusting for multiple testing, according to well-known statistical theory. This paper focuses on the estimate itself and its uncertainty, and not comparison of methods.
@@
The use and re-use of benchmark datasets has been addressed in several publications. Longjohn et al. (2025) drew attention to the role of data repositories with leaderboards and their lack of metric uncertainty. Salzberg (1997) pointed to the problem that arises because these datasets are static, and therefore run the risk of being exhausted by overfitting. Exhausted public data sets still serve important functions, for example, a new idea can show to perform satisfactory, even when it is not superior. Thompson et al. (2020) distinguished between simultaneous and sequential multiple correction procedures. In sequential testing, which is the reality of public datasets, the number of tests is constantly updated. An interesting topic they bring up is the conflict between sharing incentive, open access and stable false positive rate. There are several demonstrations of the vulnerability of re-use of benchmark datasets. Teney et al. (2020) performed a successful attack on the VQA-CP benchmark dataset (introduced by Agrawal et al., 2018); they were able to attain high performance with a non-generalisable classifier. They point towards the risk of ML methods capturing idiosyncrasies of a dataset. Torralba and Efros (2011) introduced the game Name That Dataset! motivating their concern of methods being tailored to a dataset rather than a problem. The lack of statistically significant differences among top performing methods is demonstrated by, for example, Everingham et al. (2015) and Fernández-Delgado et al. (2014).
@@
The same mechanisms of overfitting are present in public challenges through multiple submissions on the validation sample. Blum and Hardt (2015) introduced an algorithm that releases the public leaderboard score only if there has been significant improvement compared to the previous submission. This will prevent overfitting by hindering adjustments to random fluctuations in the score and not true improvements of the method. Dwork et al. (2017) used the principles of differential privacy to prevent overfitting to the validation set.
@@
Several publications show that the overfitting is not as bad as expected, and provide possible explanations. Recht et al. (2019) constructed new, independent test sets of CIFAR-10 and ImageNet. They found a dramatic drop in accuracy, around 10%. However, the order of the best performing methods is preserved, which can be explained by the fact that multiple classes (Feldman et al., 2019) and model similarity (Mania et al., 2019) slow down overfitting. Roelofs et al. (2019) compared the results from the private leaderboard (one submission per team) and the public leaderboard (several submissions per team) in Kaggle challenges. Several of their analyses indicate overfitting even though their main conclusion was that there is little evidence of overfitting.
@@
Overfitting and its consequences are avoided in most challenges by withholding the test sample so that a method can only be evaluated once and not adapted to the result. Hold-out datasets also prevents wrong use of cross-validation, see, for example, Fernández-Delgado et al. (2014) and the seminal paper of Dietterich (1998). Another contribution from public challenges is the requirement of reproducibility to collect the prize money, which is a well-known concern in science, see, for example, Gundersen and Kjensmo (2018). Public challenges typically rank the participants by a (single) performance measure, and can fall victims of Goodhart’s law, paraphrased as “When a measure becomes a target, it ceases to be a good measure.” by Teney et al. (2020). Møllersen et al. (2017) demonstrated how five different performance measures completely rearranges the leaderboard for the top participants in the ISBI 2016 challenge Skin Lesion Analysis towards Melanoma Detection (Gutman et al., 2016). Ma et al. (2021) offered a framework and platform for evaluation of NLP methods, where other aspects such as memory use and robustness are evaluated.
@@
The robustness of a method, and the variability in performance are crucial aspects that the hold-out test set of public challenges do not offer an immediate solution to. Bouthillier et al. (2021) showed empirically that hyperparameter choice and the random nature of the data are two large contributors to variance, whereas data augmentation, dropouts, and weights initialisations are minor contributors. Choice of hyperparameters are often adjusted to validation set results, and the strategies of Blum and Hardt (2015) and Dwork et al. (2017) can hinder that, if implemented by the challenge hosts.
Our contribution is a focus on better estimate for SOTA performance under the current conditions, where the approach of Bouthillier et al. (2021) does not apply. Bousquet and Elisseeff (2002) investigated how sampling randomness influences accuracy, both for regression algorithms and classification algorithms, relying on the work of Talagrand (1996) who stated that “A random variable that depends (in a “smooth” way) on the influence of many independent variables (but not too much of any of them) is essentially constant.”
@@
There is a rich literature on various aspects regarding the use of public challenges as standards for scientific publications, see, for example, Varoquaux and Cheplygina (2022) for a much broader overview than the one given here. We contribute to this conversation by shining the light on how multiplicity of classifiers’ performance on hold-out test sets create optimistic estimates of SOTA performance.
@@ 3 Multiple classifiers and biased state-of-the-art estimation
In the following sections, we take a closer look at the sample maximum estimator by presenting its distribution and investigating its bias and variance when used for SOTA estimation. We wrap the presentation in a public challenge context to refer back to previous work presented in Section 2. We use standard deviation instead of variance, because standard deviation is below $0.005$ for our numerical examples. We use the following default parameter values: $n=3,000$, $\theta=0.90$, and $m=1,000$, $\theta_{SOTA}=0.9$.
3.1 Notation and parameter values
@@
$\,\,\,\,X\sim\mathcal{B}(n,\theta)$
$\hat{\theta}(X)=\frac{X}{n}$ $:\text{estimator for }\theta$
$P(X=x)$ $:\text{probability of $x$ successes}$
$P_{x}=P(X\leq x)$ $:\text{probability of at most $x$ successes}$
$\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ $:\text{number of successes for classifiers }j=1,\ldots,m$
$C_{x}$ $:\text{number of classifiers with at most $x$ successes}$
@@
$\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$ $:\text{probability of success for classifiers }j=1,\ldots,m$
$\,\,\,\,\theta\sim P(\theta)$
$\Theta^{\prime}=\{\theta^{\prime}_{1},\theta^{\prime}_{2},\ldots,\theta^{\prime}_{m}\}$ $:\text{empirical Bayes estimate for }P(\theta)$
$\hat{\Theta}(\mathbf{X})=\{\hat{\theta}_{1}(X_{1}),\hat{\theta}_{2}(X_{2}),\ldots,\hat{\theta}_{m}(X_{3})\}$ $:\text{estimator for }\Theta$
@@
Some of the expectations, standard deviations and distributions need to be estimated by simulations. In the hierarchical set-up, we use $B=1,000$ draws from $P(\theta)$ in Eq. 1b, and $rep=100,000$ when drawing random samples of $\mathbf{X}$ for each $\theta_{j}$. We denote the estimates without the “hat”, leaning on the Law of Large Numbers and the access to arbitrary large $B$ and $rep$.
We refer to classifiers with identical probability of success as identical classifiers. We use significance level $\alpha=0.05$ and two-sided CIs.
3.2 The invention of a coin-flip classifier
We here give a motivating example, related to multiple testing, to illustrate the effect of multiplicity. For any classifier that predicts the outcome of the flip of a fair coin, the probability of success is $0.5$, and hence the SOTA performance is $\theta_{SOTA}=0.5$. If we had a public challenge where the aim was to make a coin-flip classifier, what would the SOTA performance be?
Consider an experiment where a classifier predicts the outcome of a fair coin being flipped $n=20$ times. Let the random variable $X$ denote the number of successes, $X\sim\mathcal{B}(n,\theta)$, and let ${\hat{\theta}(X)=X/n}$. We are interested in the probability of observing a large $\hat{\theta}$, say, $\geq 0.90$, and wrongfully getting the impression that a coin-flip classifier has been invented. For a single classifier, this is calculated from $1-P(X\leq x|n,\theta)$, which in our example gives us a probability of $0.00020$, reasonably low for making the false claim of coin-flip classifier invention.
For the multiplicity setting with $m$ independent classifiers, define the “success” of a classifier as $\{X\leq x\}$. Each classifier is then a Bernoulli trial. Let $C_{x}$ be the random variable that denotes the number of classifiers with success. We have that $C_{x}\sim\mathcal{B}(m,P_{x})$, where $P_{x}=P(X\leq x|n,\theta)$. Of special interest in this example is one or more classifiers performing at $0.90$ or above, defined as “failure” in this set-up. This can be written as $1-P(C_{x}=m)=1-P_{x}^{m}$. $P_{x}^{m}$ is equivalent to the probability of the sample maximum, or the $m$th order statistic, for random samples. With $m=1,0000$ classifiers, the probability of at least one classifier performing at $0.90$ or above is $0.18$, a substantial risk of grossly overestimating $\theta_{SOTA}$.
The same mechanism as described here is present when highest model performance is reported from public challenge leaderboards or public data sets. Even without the usual suspects of data re-use, overfitting, wrong use of cross-validation, multiple testing or data leakage, multiplicity, when not properly accounted for, can give the impression of the invention of a coin-flip classifier.
3.3 The probability distribution of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$
Let $\Theta=\{\theta_{1},\theta_{2},\ldots,\theta_{m}\}$ be the probability of success for $m$ classifiers, and let $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{m}\}$ be the set of random variables that denotes the number of successes (correct predictions) for each classifier on a sample of size $n$. Let the statistic $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}=\max(\{\hat{\theta}_{1}(X_{1}),\hat{\theta}_{2}(X_{2}),\ldots,\hat{\theta}_{m}(X_{m})\})$ be the naive sample maximum estimator for $\theta_{SOTA}$. We are interested in the probability distribution of $\operatorname{\hat{\theta}_{\max}(\mathbf{X})}$.
In the coin-flip example, the $\theta_{j}$’s were equal and fixed, but we can generalize to $\theta_{j}\sim P(\theta)$, where the $m$ classifiers from where $\mathbf{X}$ originates are themselves a random sample. We can model the probability mass function (pmf) as a hierarchical distribution as follows:
$\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta$ $\sim P(\operatorname{\hat{\theta}_{\max}(\mathbf{X})}|\Theta)$ (1a)
$\theta_{j}$ $\sim P(\theta).$ (1b)
The sampling procedure is illustrated in Figure 1.
Sampling procedure for the iid model